ISS Daily Summary Report – 08/21/15

H-II Transfer Vehicle (HTV)-5 Capture and Berth Readiness Status: The Space Station Remote Manipulator System (SSRMS) is in position.  All required HTV and ISS system checkouts have been successfully completed.  Capture is currently planned for ~5:55 am CDT Monday.

Human Research Program (HRP) Operations:

  • Ocular Health: Kornienko completed the ultrasound portions of his Flight Day (FD) 150 Ocular Health activities.  With Padalka’s assistance he executed a cardiac ultrasound, and with Padalka as the operator, he performed an ocular ultrasound. The ultrasound images will be used to identify changes in globe morphology, including flattening of the posterior globe, and document optic nerve sheath diameter, optic nerve sheath tortuosity, globe axial measurements, and choroidal engorgement. The Ocular Health protocol calls for a systematic gathering of physiological data to characterize the risk of microgravity-induced visual impairment/intracranial pressure in ISS crewmembers.
  • Fluid Shifts Before, During and After Prolonged Space Flight and Their Association with Intracranial Pressure and Visual Impairment (Fluid Shifts): Kelly began his FD 150 Fluid Shifts Dilution Measurements today.  Upon wakeup he collected baseline saliva, blood and urine and inserted them into the Minus Eighty Degree Celsius Laboratory Freezer for ISS (MELFI) prior to ingesting a Sodium Bromide (NaBr) tracer.  Throughout the day, he performed more urine, blood, and saliva collections and inserted those samples into MELFI as well. The Dilution Measurements are the first part in the series of FD150 Fluid Shifts measurements which investigates the causes for severe and lasting physical changes to astronauts’ eyes. Because the headward fluid shift is a hypothesized contributor to these changes, reversing this fluid shift with a lower body negative pressure device is investigated as a possible intervention. Results from this study may help to develop preventative measures against lasting changes in vision and eye damage.


Try Zero-Gravity-Asia Education Payloads Operations (EPO): Yui set up several experiments to demonstrate and promote the understanding of space and microgravity. Video of the different experiments will be downlinked and distributed. The various experiment themes were proposed by various Asian countries.


Fluid Physics Experiment Facility (FPEF) Closeout: Yui performed closeout steps following the ground software load in preparation for the Dynamic Surf-3 experiment. The investigation is part of a series of JAXA experiments addressing Marangoni convection driven by the presence of surface tension gradient as produced by a temperature difference at a liquid/gas interface. Fluid convection observations of a silicone oil liquid bridge that is generated by heating the one disc higher than the other within the Fluid Physics Experiment Facility (FPEF) are also performed. By observing and understanding how such fluids move, researchers can learn how heat is transferred in microgravity, and ultimately drive the design and development of more efficient fluid flow based systems and devices.


Extravehicular Activities (EVA) Pistol Grip Tool (PGT) Torque Analyzer Kit (TAK) Data Gather: Lindgren completed this 6-month requirement for all on-orbit PGTs. He installed the TAK on the PGTs and recorded PGT output torque data for ground analysis.


On-Board Training (OBT) Robotics On-Board Trainer (ROBoT): The USOS crew completed a session of this training during which they completed three capture point hold runs which allows free drift timing practice, malfunction response, and nominal rate approaches. They also practiced two meter approaches. This activity is in preparation for HII Transfer Vehicle (HTV)5 capture and berthing planned for next Monday, August 24.


Crew Quarters Light Replaced – Yesterday the crew swapped the dim Lamp Housing Assembly (LHA) in Crew Quarters 1 with another that had been turned off to conserve it as a spare.  12 spare LHA’s will arrive on HTV5.


Today’s Planned Activities
All activities were completed unless otherwise noted.

  • TV session with Prime Minister of the Russian Federation
  • FLUID SHIFTS – Saliva Sample Collection
  • HRF Urine Sample Collection
  • FLUID SHIFTS – Sample MELFI Insertion
  • FLUID SHIFTS – Urine Sample Collection
  • HRF – Sample MELFI Insertion
  • FLUID SHIFTS – Sample MELFI Insertion
  • FLUID SHIFTS – Blood Collection
  • FLUID SHIFTS – Blood sample collection operator
  • FLUID SHIFTS – Sample Processing
  • FLUID SHIFTS – Centrifuge Setup
  • Prep for TV PAO coverage / r/g 9678
  • Downlink of MPEG2 Video Greetings r/g 9678
  • FLUID SHIFTS – Blood Sample Transfer
  • FLUID SHIFTS – Sample MELFI Insertion
  • Camcorder Setup to View LAB RWS Monitor 3
  • Close-out ops after TV conference
  • Inventory Audit and Assessing Condition Of Main and Spare Quick Release Screw Clamps In SM r/g 9677
  • MRM2 comm config to support the P/L Ops
  • KULONOVSKYI KRISTALL. Experiment Ops. / r/g 9680
  • IMS Tagup (S-band)
  • Tropical Cyclone P/L Closeout Ops
  • FPS Tool Stowage
  • FLUID SHIFTS – Sample MELFI Insertion
  • UDOD. Experiment Ops. / r/g 9679
  • SM Ventilation System Preventive Maintenance r/g 8393
  • MRM2 Comm Reconfig for Nominal Ops / Communications System
  • KULONOVSKYI KRISTALL. Hardware Teardown / r/g 9680
  • KULONOVSKYI KRISTALL. Data Transfer to Hard Drive / r/g 9680
  • FLUID SHIFTS – Blood Collection
  • FLUID SHIFTS – Blood sample collection operator
  • FLUID SHIFTS – Centrifuge Setup
  • FLUID SHIFTS – Sample MELFI Insertion
  • FPS Tool Stowage
  • RWS Powerup
  • FLUID SHIFTS – Blood Sample Transfer
  • FLUID SHIFTS – Sample MELFI Insertion
  • TRY ZERO G Experiment Ops
  • CARDIOLAB HEMATOCRIT Centrifuge Checkout
  • Swap condensation collection from the Lab Condensate Tank to the WPA WW tank while the Sabatier Jumper is installed
  • WRM. Condensate Pumping Initiation
  • FS – Saliva Sample Collection
  • FLUID SHIFTS – Sample MELFI Insertion
  • FLUID SHIFTS – Urine Sample Collection
  • USND2 – Activation of Ultrasound Hardware and Camcorder
  • OH-CARDIAC – Preparation
  • FLUID SHIFTS – Sample MELFI Insertion
  • Ocular Health – Cardiac and Blood Pressure Operations (Subject)
  • WRM. Condensate Pumping Terminate
  • Ocular Health – Cardiac and Blood Pressure Operations (Operator)
  • CONTUR-2. Hardware Installation on panel 418
  • FPEF – Connector Removal
  • Ocular Health – Closeout Ops
  • CONTUR-2. Experiment Session Ops. / r/g 9675
  • Ocular Health – Data Export
  • GREAT START. Preparation for the Experiment Ops / r/g 9671
  • OBT – HTV Capture OBT
  • GREAT START. Video of the Experiment Ops / r/g 9671
  • SPHERES – Hardware Stow after the experiment
  • Transfer of IVA Tissue-Equivalent Proportional Counter (IV-TEPC)
  • On-orbit Hearing Assessment using EARQ
  • GREAT START. Closeout ops / r/g 9671
  • Ultrasound – Scanning Prep
  • HMS – Ultrasound2 – Scan by CMO
  • Ultrasound 2 – Scan
  • CONTUR-2. Experiment Session Ops.r/g 9675
  • Ultrasound Data Export
  • HTV Transfer List Review
  • CONTUR-2. Hardware Disassembly from panel 418
  • HMS – Ultrasound2 – Scan by CMO
  • HTV Transfer List Conference
  • Ultrasound 2 – Scan
  • Retrieve NOD1SD2 Lamp Housing Assembly (LHA) for installation in Deck Crew Quarters [Completed yesterday]
  • HMS – Food Frequency Questionnaire
  • SHD – Weekly Questionnaire
  • Ultrasound Data Export
  • Deck Crew Quarters LHA (Lamp Housing Assembly) R&R [Completed yesterday]
  • SHD – Weekly Questionnaire
  • HMS – Ultrasound 2 System – Scan Closeout Ops
  • Video Downlink End
  • Install Lamp Housing Assembly (LHA) removed from Deck Crew Quarters at NOD1SD2 [Completed yesterday]
  • CONTENT. Experiment Ops / r/g 9676
  • USND2 – Hardware Deactivation and Closeout Ops


Completed Task List Items

  • None


Ground Activities

All activities were completed unless otherwise noted.

  • ROBoT training support
  • Cardiac/Ultrasound ops
  • MSS powerup
  • Ground Camera Calibration Part 2


Three-Day Look Ahead:

Saturday, 08/22: Crew off duty, Housekeeping

Sunday, 08/23: Crew off duty, HTV Hardware Command Panel setup

Monday, 08/24: HTV5 capture/berthing


QUICK ISS Status – Environmental Control Group:

                              Component Status
Elektron On
Vozdukh Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) Off
[СКВ] 2 – SM Air Conditioner System (“SKV2”) On
Carbon Dioxide Removal Assembly (CDRA) Lab Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 Operate
Major Constituent Analyzer (MCA) Lab Shutdown
Major Constituent Analyzer (MCA) Node 3 Operate
Oxygen Generation Assembly (OGA) Process
Urine Processing Assembly (UPA) Standby
Trace Contaminant Control System (TCCS) Lab Off
Trace Contaminant Control System (TCCS) Node 3 Full Up